Status Report

NOAA SATOPS Morning Report: Friday, May 25, 2012

By SpaceRef Editor
May 25, 2012
Filed under ,

SATOPS Morning Report: Friday, May 25, 2012

Customers or users who want to call to ask about problems or report outages can use the following contacts:
– For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-763-8444
– To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: [email protected]

DOY 146

GOES
GOES-12 Imager instrument Cycle slips caused intermittent image degradation at DOY145 9:45z,10:45z,11:45 DOY146 02:28z, 02:45z, 03:28z & 0345Z

All other GOES operations were nominal over the past 24 hours.
No anomalies affecting product delivery via GVAR or s/c health & safety.
For GOES status and other information please see:
http://www.oso.noaa.gov/goes

POES
All POES operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For a detailed status and description of POES see:
http://www.oso.noaa.gov/poes

DMSP
Issues with POGO-B resolved.

All other DMSP operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For more DMSP Information please see:
http://dmsp.ngdc.noaa.gov/dmsp.html

JASON-2
All other JASON-2 operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.

ESPC Reported data outage and delays
1830-1915z GOES-15 Missed GINI AWIPS Images due to high ber at wcda, Product(s) or Data Impacted: 1830Z (HAWNATL,WCONUS). 1840Z/1910Z (WCONUS), 1845Z /1915Z (AKNATL, AKREG, HAWNATL, HAWREG, WCONUS), 1900Z (AKNATL, AKREG, HAWNATL, HAWREG, WCONUS, NHEMI AND SUPERN)

Questions or comments to [email protected]

SpaceRef staff editor.